# Towards Unified Latent Space for 3D Molecular Latent Diffusion Modeling

### v2

@@ 4.1 Experimental Setup
Datasets. Following previous studies Hoogeboom et al. (2022); Huang et al. (2023b); Vignac et al. (2023), we conduct experiments on the popular QM9 Ramakrishnan et al. (2014) and the GEOM-Drugs Axelrod and Gómez-Bombarelli (2020) datasets.
  • QM9 Ramakrishnan et al. (2014): Contains 130k small organic molecules (up to 9 heavy atoms, 29 total atoms) with quantum chemical properties. Following standard practice Hoogeboom et al. (2022); Huang et al. (2023b), we use 100K/18K/13K train/val/test splits.
  • GEOM-Drugs Axelrod and Gómez-Bombarelli (2020): Pharmaceutical-scale dataset with 450k drug-like molecules (average 44.4 atoms, max 181 atoms). We use 80%/10%/10% splits and retain the lowest-energy conformer per molecule.
Baselines. For de novo 3D molecule generation, we compare UDM-3D with CDGS Huang et al. (2023a), JODO Huang et al. (2023b), MiDi Vignac et al. (2023), G-SchNet Gebauer et al. (2019), G-SphereNet Luo and Ji (2022), EDM Hoogeboom et al. (2022), MDM Huang et al. (2023c), GeoLDM Xu et al. (2023). For conditional 3D molecule generation, we compare UDM-3D with EDM Hoogeboom et al. (2022), EEGSDE Bao et al. (2023), GeoLDM Xu et al. (2023), and JODO Huang et al. (2023b).
@@ 4.2 De Novo 3D Molecule Generation
Experimental Setting. Generating a 3D molecule involves generating the structural validity (atom/bond features) and accurate spatial arrangements (3D coordinates). Therefore we adapt the comprehensive metrics from Huang et al. (2023b); Hoogeboom et al. (2022):
  • 2D Metrics: Atom stability, validity & completeness (V&C), validity & uniqueness (V&U), validity & uniqueness & novelty (V&U&N), similarity to nearest neighbor (SNN), fragment similarity (Frag), scaffold similarity (Scaf), and Fréchet ChemNet Distance (FCD) Polykovskiy et al. (2020).
  • 3D Metrics: Atom stability, FCD, maximum mean discrepancy (MMD) Gretton et al. (2012) for the distributions of bond lengths, bond angles, and dihedral angles.
The details of these metrics are in Appendix A.2.1. We also report all the metrics for molecules from training set, serving as an approximate upper bound for performance.
@@
  1. 1.
    Breakthrough in Geometric Accuracy. Our unified latent space proves highly effective for 3D geometric modeling which is crucial for accurate 3D molecule generation. UDM-3D achieves significantly better performance than baselines for the distributional distance of bond lengths, bond angles, reducing errors by an order of magnitude. This demonstrates UDM-3D’s ability to generate highly accurate 3D structures. For the distributional distance of dihedral angles, a key measurement for 3D conformation accuracy Jing et al. (2022), UDM-3D also achieves the best performance. Notably, these advantages on stereochemical geometric structures persist for complex drug-like molecules in GEOM-Drugs, where UDM-3D improves the distributional distance for bond length by 25 times compared to GeoLDM (9.89E-03 v.s. 3.91E-01). This quantum leap stems from UDM-3D’s advantage to jointly optimize chemical and 3D geometric constraints when generating within a unified latent space, ensuring consistency across different molecular modalities.
  2. 2.
    Novel and High Quality Generation. UDM-3D achieves remarkable V&U&N scores of 0.948 (QM9) and 0.958 (GEOM-Drugs) while maintaining stability on par with top baselines like JODO. This demonstrates UDM-3D’s unique capability to generate novel molecules while preserving high-quality structures. The improvement is attributed to the unified latent space: by encoding atom types and bond features in a unified space, the model avoids overfitting to common molecular patterns while maintaining chemical validity.
@@ 4.3 Conditional 3D Molecule Generation
Experimental Settings. We evaluate conditional generation of molecules with target quantum properties using the protocol from Hoogeboom et al. (2022); Xu et al. (2023). Specifically, our target properties include $C_{v}$, $\mu$, $\alpha$, $\epsilon_{\text{HOMO}}$, $\epsilon_{\text{LUMO}}$, $\Delta\epsilon$, and we evaluate the Mean Absolute Error (MAE) between target and predicted properties. More details on the properties and settings are provided in Appendix A.2.2.
Table 4 presents the MAE results for conditional generation on QM9. UDM-3D achieves the lowest MAE across five properties, demonstrating its effectiveness in generating molecules with desired target properties. Notably, UDM-3D reduces the average MAE by 42.2% compared to GeoLDM, with strong gains (52.7% relative improvement) in $\Delta\epsilon$ (HOMO-LUMO gap), demonstrating significant improvements in modeling quantum chemical properties. These gains highlight the critical role of our unified latent space in harmonizing multi-modal information during conditional generation. While GeoLDM struggles to correlate all the modalities with target quantum properties due to its fragmented latent spaces, our unified latent representation inherently preserves the interplay between 3D geometry, bonds, and electronic characteristics, enabling precise modeling over molecule-property relationships.
@@ 4.4 Training & Sampling Efficiency
(a) Training time comparison.
(b) Sampling speed comparison.
Figure 3: Efficiency comparison. (a): Our UDM-3D adopts a two-stage pipeline (stacked bar: 14h VAE training + 38h LDM training), while baselines use end-to-end training.
@@
Training Efficiency. As shown in Figure 3(a), UDM-3D completes training in 52 hours (14h UAE-3D + 38h UDM-3D), which is 5.3 times faster than GeoLDM (449h) and 2.7 times faster than JODO (139h). This efficiency stems from: (1) Our decoupled training paradigm: UAE-3D first learns 3D molecule compression, allowing the DiT to focus solely on generative modeling of the compressed latents. (2) Efficient diffusion with DiT: Unlike previous molecular diffusion models that require complex architectures for multi-modality and equivariance, DiT’s simple and highly parallelizable design allows for faster training and sampling on the unified latent space. Crucially, our speedup does not compromise geometric accuracy, as guaranteed by UAE-3D’s near-lossless molecular reconstruction (cf. Table 1).
@@
Sampling Speed. As shown in Figure 3(b), UDM-3D generates molecules in just 0.081s per sample, making it 7.3x faster than EDM/GeoLDM (0.59s) and 9.8x faster than JODO (0.79s). This speed advantage comes from the unified latent space, which simplifies DiT’s modeling and avoids complex neural architectures.
4.5 Ablation Studys
4.5.1 DiT Backbone
We conduct an ablation study on DiT by comparing it to a vanilla Transformer for de novo 3D molecule generation on the QM9 dataset (Table 5). The results show that DiT outperforms the vanilla Transformer across all metrics, which we attribute to its adaptive LayerNorm layers. These layers enable DiT to effectively handle data with varying noise scales, improving diffusion performance.
@@ 4.5.1 DiT Backbone
Table 5: Comparing DiT with vanilla Transformer for latent diffusion modeling. Models have the same depth and hidden size. A.S. stands for atom stability
Arch. A.S.${}_{\text{3D}}$ A.S.${}_{\text{2D}}$ V&C V&U V&U&N
DiT 0.993 0.999 0.983 0.972 0.948
Trans. 0.983 0.997 0.938 0.922 0.922
4.5.2 Bonded Distance Loss
Table 6 presents an ablation study on the weight $\gamma_{d}$ for the bonded distance loss on the QM9 dataset. We vary $\gamma_{d}$ to analyze its effect on reconstruction accuracy and the preservation of bond geometries. The results indicate that setting $\gamma_{d}=10$ achieves an optimal balance between compressing coordinate information and maintaining geometric fidelity. This experiment validates the necessity of tuning the bonded distance loss weight to enhance model performance.
@@ 4.5.2 Bonded Distance Loss
Table 6: The influence of different bonded distance loss weight $\gamma_{d}$ on the QM9 dataset.
$\gamma_{d}$ / VAE Rec. Atom Accuracy (%) Bond Accuracy (%) Coordinate RMSD (Å)
0 (w/o $\gamma_{d}$) 1.000 1.000 0.0034
@@
1 0.996 0.959 0.962 0.940 0.940 0.494 0.978 0.874 0.283
5 0.997 0.965 0.970 0.947 0.947 0.500 0.983 0.888 0.255
10 0.999 0.988 0.983 0.972 0.948 0.508 0.987 0.897 0.161
20 0.999 0.988 0.980 0.969 0.947 0.507 0.987 0.898 0.172
$\gamma_{d}$ / 3D-Metric AtomStable MolStable Bond $\downarrow$ Bond $\downarrow$ Dihedral $\downarrow$
1 0.985 0.857 1.25E-01 1.80E-02 1.06E-03
5 0.989 0.909 1.12E-01 1.19E-02 7.18E-04
10 0.993 0.929 7.04E-02 9.84E-03 3.47E-04
20 0.993 0.929 7.05E-02 9.84E-03 3.45E-04
4.5.3 SE(3)-Equivariant Augmentations
Table 7 shows the influence of various SE(3)-equivariant augmentations strategies on the performance over the QM9 dataset. The results reveal that data augmentation improves overall performance by providing diverse geometric views during UDM-3D training.
In addition to the standard performance metrics, we further investigate the reconstruction error of UAE-3D under different 3D transformations. Table 8 presents the Root Mean Square Deviation (RMSD) of the reconstructed molecules.
Table 7: The effect of SE(3)-equivariant augmentations strategies on the QM9 dataset.
[…]
w/o aug. 0.995 0.947 0.950 0.935 0.921 0.492 0.971 0.799 0.581
Rot. 0.995 0.948 0.951 0.943 0.927 0.493 0.979 0.875 0.315
Trans. 0.999 0.986 0.980 0.967 0.944 0.507 0.981 0.884 0.202
Trans. + Rot. 0.999 0.988 0.980 0.969 0.947 0.507 0.987 0.898 0.172
3D-Metric AtomStable MolStable Bond $\downarrow$ Bond $\downarrow$ Dihedral $\downarrow$
w/o aug. 0.993 0.879 1.01E-01 1.18E-02 1.08E-03
Rot. 0.993 0.876 8.12E-02 9.99E-03 3.90E-04
Trans. 0.989 0.896 7.08E-02 1.04E-02 6.81E-04
Trans. + Rot. 0.993 0.929 7.05E-02 9.84E-03 3.45E-04
Table 8: Reconstruction error analysis under different SE(3)-equivariant augmentations strategies on the QM9 dataset.
Augmentation RMSD
w/o aug. 0.0002
Rotation 0.0065
Translation 0.0048
Rotation + Translation 0.0067
4.5.4 Latent Dimensions
@@
To investigate the impact of latent dimensionality on 3D molecule generation, we conduct an ablation study on the QM9 dataset. Table 9 summarizes the performance metrics across different latent space dimensions (4, 8, 16, and 32). The results indicate that a too-low latent dimension (e.g., 4) fails to capture the necessary information, leading to inferior performance. Increasing the dimension to 16 results in substantial performance improvements, while further increasing to 32 does not yield significant additional gains. This demonstrates that a latent dimension of 16 provides an optimal trade-off between model capacity and reconstruction fidelity.
[…]
4 0.914 0.325 0.580 0.566 0.566 3.27E-01 2.20E-01 4.19E-03
8 0.987 0.882 0.942 0.923 0.923 1.29E-01 1.77E-02 8.14E-04
16 0.999 0.988 0.980 0.969 0.947 7.05E-02 9.84E-03 3.45E-04
32 0.986 0.867 0.943 0.918 0.918 9.98E-02 1.42E-02 7.77E-04
5 Related Works
3D Molecule Generation. Early efforts for 3D molecule generation focus on autoregressive approaches Gebauer et al. (2019, 2021); Luo and Ji (2022), constructing 3D molecules sequentially by adding atoms or molecular fragments. With the success of diffusion models across various domains Sohl-Dickstein et al. (2015); Ho et al. (2020); Dhariwal and Nichol (2021), they also become the de facto method for 3D molecule generation Hoogeboom et al. (2022); Bao et al. (2023). However, the early diffusion works can easily generate invalid molecules because of overlooking the bond information. To bridge this gap, subsequent works Huang et al. (2023b); Vignac et al. (2023); Hua et al. (2023) additionally consider bond generation by building a separate diffusion process. However, generating different modalities in separate diffusion processes unnecessarily complicates the model design, reducing efficiency. It also risks compromising the consistency between these modalities. To address this, our UDM-3D performs generative modeling in UAE-3D’s unified latent space that integrates all molecular modalities, improving efficiency and generation quality.
Latent Diffusion Models (LDMs) for 3D Molecule Generation. To improve the efficiency, LDMs employ a VAE to compress raw data into a lower-dimensional latent space, where a diffusion model performs generative modeling Rombach et al. (2022). This approach has been very popular for image generation Peebles and Xie (2023). However, existing LDM-based methods for 3D molecule generation still face challenges in separated latent spaces. For example, GeoLDM Xu et al. (2023) compresses atomic features and 3D coordinates separately, failing to achieve a unified latent space distribution. In 3D protein molecule generation, PepGLAD Kong et al. (2024) compresses 1D sequences and 3D structures into two separate latent spaces. Our UDM-3D addresses these issues by employing a unified latent space, significantly improved generation efficiency.

### v3

@@ 4.1 Experimental Setup
Datasets. Following prior works [5, 6, 9], we conduct experiments on QM9 [24] and GEOM-Drugs [25] datasets. QM9 [24] contains 130k small organic molecules (up to 9 heavy atoms) with quantum chemical properties. We use 100K/18K/13K train/val/test splits, following [5, 6]. GEOM-Drugs [25] is a pharmaceutical-scale dataset with 450k drug-like molecules (average 44.4 atoms, max 181 atoms). We use 80%/10%/10% splits and retain the lowest-energy conformer per molecule.
Baselines. For de novo 3D molecule generation, we compare UDM-3D with CDGS [32], JODO [6], MiDi [9], G-SchNet [12], G-SphereNet [33], EDM [5], MDM [34], GeoLDM [17], and ADiT [35]. For conditional generation, we additionally use baselines of EEGSDE [8] and GeoBFN [36]. We do not report MDM’s performance on GEOM-DRUGS because we are unable to reproduce reasonable results using their official code. Similarly, ADiT’s performance on GEOM-DRUGS is omitted, because it is not available in their paper, code, and released checkpoint.
@@ 4.2 De Novo 3D Molecule Generation
Experimental Setting. Generating a 3D molecule involves generating the structural validity (atom/bond features) and accurate spatial arrangements (3D coordinates). Therefore, we adapt the comprehensive metrics from [6, 5]. We also report all the metrics for molecules from training set, serving as an approximate upper bound for performance. Our evaluation metrics can be divided into two groups: (1) 2D Metrics: Atom stability, validity & completeness (V&C), validity & uniqueness (V&U), validity & uniqueness & novelty (V&U&N), similarity to nearest neighbor (SNN), fragment similarity (Frag), scaffold similarity (Scaf), and Fréchet ChemNet Distance (FCD) [3]; (2) 3D Metrics: Atom stability, FCD, maximum mean discrepancy (MMD) [37] for the distributions of bond lengths, bond angles, and dihedral angles. More experimental details are in Appendix D.1.
Table 2: Performance of de novo 3D molecule generation on QM9. * indicates results reproduced using official source codes, while other baseline results are taken from [6]. ADiT’s relatively poor performance stems from differences in evaluation protocols -— their generated molecules exhibit significant stability issues. We provide a detailed analysis in Appendix D
[…]
Train 0.063 0.999 0.988 0.989 0.989 0.000 0.490 0.992 0.946
CDGS 0.798 0.997 0.951 0.951 0.936 0.860* 0.493 0.973 0.784
JODO 0.138 0.999 0.988 0.990 0.960 0.780* 0.522 0.986 0.934
MiDi* 0.187 0.998 0.976 0.980 0.954 0.769 0.501 0.979 0.882
EQGAT-diff* 2.088 0.999 0.971 0.965 0.950 0.891 0.482 0.950 0.703
UDM-3D, ours ${\textbf{0.130}}^{\color[rgb]{1,.5,0}-5.80\%}$ 0.999 0.988 0.983 0.973 0.950 0.508 0.987 0.898
${}_{\text{3D}}$$\downarrow$ AtomStable MolStable Bond $\downarrow$ Bond $\downarrow$ Dihedral $\downarrow$
Train 0.877 0.994 0.953 5.44E-04 4.65E-04 1.78E-04
EDM 1.285 0.986 0.817 1.30E-01 1.82E-02 6.64E-04
MDM 4.861 0.992 0.896 2.74E-01 6.60E-02 2.39E-02
JODO 0.885 0.992 0.934 1.48E-01 1.21E-02 6.29E-04
GeoLDM 1.030 0.989 0.897 2.40E-01 1.00E-02 6.59E-04
MiDi* 1.100 0.983 0.842 8.96E-01 2.08E-02 8.14E-04
EQGAT-diff* 1.520 0.988 0.888 4.21E-01 1.89E-02 1.24E-03
ADiT* 2.884 0.211 0.000 9.98E-01 3.38E-02 1.46E-03
[…]
@@
Breakthrough in Geometric Accuracy. Our unified latent space enables highly accurate 3D geometric modeling, a key factor in realistic molecule generation. UDM-3D significantly outperforms baselines in distributional distances of bond lengths and angles, reducing errors by an order of magnitude. These improvements hold even for complex, drug-like molecules in GEOM-Drugs—where UDM-3D achieves a 25x lower bond length error than GeoLDM (9.89E-03 v.s. 3.91E-01). This leap in accuracy comes from UDM-3D’s ability to jointly optimize chemical and geometric constraints in a unified latent space, maintaining consistency across molecular modalities.
Novel and High Quality Generation. UDM-3D achieves outstanding V&U&N scores of 0.907 on GEOM-Drugs and 0.950 on QM9, while achieving state-of-the-art 3D stability and comparable 2D stability to strong baselines such as JODO. This highlights its ability to generate novel molecules without compromising structural quality. The key lies in the unified latent space, which jointly encodes atom types, bond types, and coordinates, enabling the model to capture complex interdependencies among molecular features.
@@ 4.3 Conditional 3D Molecule Generation
Experimental Settings. We evaluate conditional generation of molecules with target quantum properties using the protocol from [5, 17]. Specifically, our target properties include $C_{v}$, $\mu$, $\alpha$, $\epsilon_{\text{HOMO}}$, $\epsilon_{\text{LUMO}}$, $\Delta\epsilon$, and we evaluate the Mean Absolute Error (MAE) between target and predicted properties. More details on the properties and settings are provided in Appendix D.2.
Table 3 reports MAE results for conditional generation on QM9. UDM-3D achieves the lowest MAE in five out of six properties, showing strong capability in generating molecules that match target properties. On average, it reduces MAE by 42.2% compared to GeoLDM, with a notable 52.7% improvement in predicting the HOMO-LUMO gap ($\Delta\epsilon$). These results highlight the importance of our unified latent space that effectively integrates all molecular modalities during generation. While GeoLDM struggles to correlate all the modalities due to its fragmented latent spaces, our unified latent representation inherently preserves the interplay between 3D geometry, bonds, and electronic characteristics, enabling precise modeling over molecule-property relationships.
Table 3: Mean Absolute Error (MAE) for conditional 3D molecule generation on QM9.
U-Bound 1.613 8.98 6.879 645 1457 1464
L-Bound 0.043 0.09 0.040 039 0036 0065
EDM 1.123 2.78 1.065 371 601 671
EEGSDE 0.777 2.50 0.941 302 447 487
GeoLDM 1.108 2.37 1.025 340 522 587
GeoBFN 0.998 2.34 0.949 328 516 577
JODO 0.628 1.42 0.581 226 256 335
[…]
@@ 4.4 Training & Sampling Efficiency
Training Efficiency. As Figure 2(c) shows, UDM-3D’s training costs 52 hours (14h UAE-3D + 38h UDM-3D), which is 5.3 times faster than GeoLDM and 2.7 times faster than JODO. This efficiency stems from: (1) Our decoupled training paradigm: UAE-3D first learns 3D molecule compression, allowing DiT to focus solely on generative modeling of the compressed latents. (2) Efficient diffusion with DiT: Unlike previous molecular diffusion models that require complex architectures for multi-modality and equivariance, DiT’s simple and highly parallelizable design allows for faster training and sampling in the unified latent space. Crucially, our speedup does not compromise performance, as guaranteed by UAE-3D’s near-lossless molecular reconstruction (cf. Figure 2(a;b)).
@@
Sampling Speed. Figure 2(d) shows that UDM-3D generates each molecule in just 0.081s, which is 7.3x faster than EDM/GeoLDM and 9.8x faster than JODO. This speed advantage comes from our unified latent space, which simplifies DiT’s modeling and avoids complex neural architectures.
(a) Translation.
(b) Rotation.
(c) Rot. $\rightarrow$ Trans.
Figure 4: t-SNE visualizations of UAE-3D’s latents under SE(3) augmentations. (a) Translations along a fixed direction. (b) Rotations along a fixed axis. (c) Sequential rotations followed by translations. Color gradients show increasing distances or angles.
@@
Table 4: Comparing DiT with Transformer for latent diffusion modeling. Models have the same depth and hidden size. A.S. $\rightarrow$ atom stability.
Metric DiT Transformer
A.S.${}_{\text{3D}}$ 0.993 0.983
A.S.${}_{\text{2D}}$ 0.999 0.997
V&C 0.983 0.938
V&U 0.973 0.922
V&U&N 0.950 0.922
Table 5: The performance influence of SE(3)-equivariant augmentations on the QM9 dataset.
[…]
No aug. 0.581 0.995 0.947 0.950 0.935 0.921 0.492 0.971 0.799
+ Rot. 0.315 0.995 0.948 0.951 0.943 0.927 0.493 0.979 0.875
+ Trans. 0.202 0.999 0.986 0.980 0.967 0.944 0.507 0.981 0.884
+ Trans. + Rot. 0.130 0.999 0.988 0.983 0.973 0.950 0.508 0.987 0.898
${}_{\text{3D}}$$\downarrow$ AtomStable MolStable Bond $\downarrow$ Bond $\downarrow$ Dihedral $\downarrow$
No aug. 1.065 0.993 0.879 1.01E-01 1.18E-02 1.08E-03
+ Rot. 1.007 0.993 0.876 8.12E-02 9.99E-03 3.90E-04
+ Trans. 0.902 0.989 0.896 7.08E-02 1.04E-02 6.81E-04
+ Trans. + Rot. 0.881 0.993 0.935 7.04E-02 9.84E-03 3.47E-04
4.5 Analysis and Ablation Studies
UAE-3D’s Latents Present Structured Variations across Geometric Movements. Figure 4 shows the t-SNE visualizations of UAE-3D’s latent representation under different SE(3) augmentations. Each dot represents a molecular embedding in the latent space, and the color gradient from light to dark corresponds to an increase in the translation distance (middle) or the rotation angle (right) for the same molecule. We observe that the same molecule’s representations remain close in the latent space and as the augmentation scale increases, the representation changes along a consistent direction. Figure 4(c) shows that the representations change moving directions when the augmentation is switched from rotation to translation. These observations demonstrate that UAE-3D captures meaningful and structured variations in molecular representations across SE(3) augmentations.
3D Molecule Samples by UDM-3D. We present 3D molecules generated by our method as case studies in Figure 5. The generated molecules are chemically valid and exhibit diverse and complex structures, showcasing UDM-3D’s ability to generate realistic 3D molecular conformations.
Ablating the DiT Backbone. We compare DiT with a vanilla Transformer for de novo 3D molecule generation on the QM9 dataset (cf. Table 4). The results show that DiT outperforms the vanilla Transformer across all metrics, which we attribute to its adaptive LayerNorm layers. These layers enable DiT to effectively handle data with varying noise scales, improving diffusion performance.
SE(3)-Equivariant Augmentations. Table 5 shows the performance influence of various SE(3)-equivariant augmentations on the QM9 dataset. The results reveal that data augmentation improves overall performance by providing diverse geometric views during UDM-3D training.
In addition to the standard performance metrics, we further investigate UAE-3D’s reconstruction error on the test set when trained under different 3D transformations. Table 6 presents the RMSD of the reconstructed molecules. By including both rotation and translation augmentations, we observe a significant reduction in reconstruction error (from $1.4\times 10^{-4}$ to $0.2\times 10^{-4}$), achieving the lowest RMSD. This shows that our SE(3)-equivariant augmentations effectively enhance the model’s ability to learn geometric consistency by exploring a more comprehensive 3D molecular space.
Table 6: RMSD results when trained with different SE(3) augmentations on QM9.
Train with RMSD (×$\times 10^{-4}$Å)
No aug. 1.4
+ Rot. 1.1
+ Trans. 0.9
+ Trans. + Rot. 0.2
Figure 5: Visualization of random samples generated by UDM-3D on QM9.
5 Related Works
3D Molecule Generation. Early efforts for 3D molecule generation focus on autoregressive approaches [12, 13, 33], constructing 3D molecules sequentially by adding atoms or molecular fragments. With the success of diffusion models across various domains [38, 15, 39], they also become the de facto method for 3D molecule generation [5, 8]. However, the early diffusion works can easily generate invalid molecules because of overlooking the bond information. To bridge this gap, subsequent works [6, 9, 11] additionally consider bond generation by building a separate diffusion process. However, generating different modalities in separate diffusion processes unnecessarily complicates the model design, reducing efficiency. It also risks compromising the consistency between these modalities. To address this, our UDM-3D performs generative modeling in UAE-3D’s unified latent space that integrates all molecular modalities, improving efficiency and generation quality.
LDMs for 3D Molecule Generation. For efficiency, LDMs employ a VAE [16] to compress raw data into a low-dimensional latent space, where a diffusion model performs generative modeling [14]. This method has been very popular for image generation [23]. However, existing 3D molecular LDMs still face challenges in separated latent spaces. For example, GeoLDM [17] compresses atom features and 3D coordinates separately, failing to build a unified latent space. Other works [36, 40] process different molecular components (atoms/subgraphs or atom types/coordinates) in separate channels, increasing modeling complexity. In 3D protein generation, PepGLAD [18] compresses 1D sequences and 3D structures into two separate latent spaces. Our UDM-3D addresses these issues by employing a unified latent space, significantly improved generation efficiency.
6 Conclusion and Future Works
In this paper, we propose UAE-3D to compress the multi-modal features of 3D molecules into a unified latent space, and demonstrate the effectiveness of latent diffusion modeling on this space by introducing UDM-3D. By integrating atom types, chemical bonds, and 3D coordinates into a single latent space, our model effectively addresses the inherent challenges of multi-modality and SE(3) equivariance for 3D molecule generation. Extensive experiments on GEOM-Drugs and QM9 confirm leading performances in both de novo and conditional 3D molecule generation, setting new benchmarks for both quality and efficiency. UAE-3D’s success highlights the benefits of building a unified latent space for molecular design. Moving forward, we will transfer this unified latent space to new modalities, including proteins and RNAs, and extend it to broader molecular design tasks, including structure-based drug design.
Appendix A Limitations and More Future Works
Scope. In this work, we explore 3D molecule generation under the settings of unconditional generation and conditional generation targetting at quantum chemical properties. Other molecule generation tasks, such as structure-based drug design [41], is out of the scope of this work. We leave this exploration to future works.
Molecule Size. Our method relies on a separate module to decide each generated molecule’s number of atom, following prior works [6, 5, 8]. Specifically, we use the molecule size distribution measured on the training dataset to sample new molecule size. While this method shows decent performance on the tested tasks, we conjecture this will be one of the bottlenecks when the tested task become more complex. Resolving this issue might demand the incorporation of auto-regressive based method for molecule generation, in which the generated molecule size is automatically controlled by the auto-regressive process.
Future Works: Multi-modal Molecule Language Models. While UAE-3D’s latents are near-lossless compressions of the original molecule, it inspires the exploration of jointly modeling UAE-3D’s latents with text sequences for text-guided 3D molecule generation, following [42, 43, 44]
Appendix B More Experimental Results
B.1 More Ablation Studies and Analysis
[…]
4 0.914 0.325 0.580 0.566 0.566 3.27E-01 2.20E-01 4.19E-03
8 0.987 0.882 0.942 0.923 0.923 1.29E-01 1.77E-02 8.14E-04
16 0.999 0.988 0.983 0.973 0.950 7.04E-02 9.84E-03 3.47E-04
32 0.986 0.867 0.943 0.918 0.918 9.98E-02 1.42E-02 7.77E-04
@@ B.1 More Ablation Studies and Analysis
Table 8: The influence of different bonded distance loss weight $\lambda$ on the QM9 dataset.
$\lambda$ / VAE Recon. Atom Accuracy (%) Bond Accuracy (%) Coordinate RMSD (Å)
0 (w/o $\lambda$) 1.000 1.000 0.0034
@@
1 0.283 0.996 0.959 0.962 0.940 0.940 0.494 0.978 0.874
5 0.255 0.997 0.965 0.970 0.947 0.947 0.500 0.983 0.888
10 0.130 0.999 0.988 0.983 0.973 0.950 0.508 0.987 0.898
20 0.172 0.999 0.988 0.980 0.969 0.947 0.507 0.987 0.898
$\lambda$ / ${}_{\text{3D}}$$\downarrow$ AtomStable MolStable Bond $\downarrow$ Bond $\downarrow$ Dihedral $\downarrow$
1 0.901 0.985 0.857 1.25E-01 1.80E-02 1.06E-03
5 0.892 0.989 0.909 1.12E-01 1.19E-02 7.18E-04
10 0.881 0.993 0.935 7.04E-02 9.84E-03 3.47E-04
20 0.889 0.993 0.929 7.05E-02 9.84E-03 3.45E-04
Table 9: Reconstruction accuracy (%) for atom and bond types, and coordinate RMSD on the test sets.
Dataset Method Atom Accuracy (%) Bond Accuracy (%) Coordinate RMSD (Å)
GEOM-Drugs GeoLDM 096.9 093.7 0.2526
UDM-3D, ours 100.0 100.0 0.0008
QM9 GeoLDM 098.6 096.2 0.1830
ADiT 085.7 - 0.3598
UDM-3D, ours 100.0 100.0 0.0002
Reconstruction Errors. To assess the reconstruction fidelity of UAE-3D, we evaluate its performance on the held-out test sets of QM9 and GEOM-Drugs. We report atom-type and bond-type accuracies as well as the coordinates RMSD, which jointly reflect the VAE’s ability to compress and reconstruct both invariant and equivariant molecular features. Table 9 presents the reconstruction results. On the QM9 dataset, UDM-3D achieves perfect reconstruction accuracy (100% for both atoms and bonds) with a near-zero coordinate RMSD of 0.0002 Å, substantially overperforming GeoLDM and ADiT. Notably, ADiT does not retain and reconstruct chemical bond information, which is a limitation when reconstructing 3D molecules. On the more complex GEOM-Drugs dataset, UDM-3D again achieves perfect reconstruction (100.0% atom and bond accuracy, 0.0008 Å RMSD), whereas GeoLDM exhibits noticeable degradation. This demonstrates that UDM-3D’s unified latent space not only faithfully compresses all molecular modalities but also generalizes robustly to large and diverse molecular datasets. These results show that UAE-3D effectively learns a unified latent representation that preserves both the chemical identity and geometric precision of molecules. These findings align with the visual comparison in Figure 2(a;b), and reinforce our claim that UAE-3D achieves near-lossless reconstruction—an essential prerequisite for accurate latent diffusion modeling.
@@
Influenece of Latent Dimensions. To investigate the impact of latent dimensionality on 3D molecule generation, we conduct an ablation study on the QM9 dataset. Table 7 summarizes the performance metrics across different latent space dimensions (4, 8, 16, and 32). The results indicate that a too-low latent dimension (e.g., 4) fails to capture the necessary information, leading to inferior performance. Increasing the dimension to 16 results in substantial performance improvements, while further increasing to 32 does not yield significant additional gains. This demonstrates that a latent dimension of 16 provides an optimal trade-off between model capacity and reconstruction fidelity.
Bonded Distance Loss. Table 8 presents an ablation study on the weight $\lambda$ for the bonded distance loss on the QM9 dataset. We vary $\lambda$ to analyze its effect on reconstruction accuracy and the preservation of bond geometries. The results indicate that setting $\lambda=10$ achieves an optimal balance between compressing coordinate information and maintaining geometric fidelity. This experiment validates the necessity of tuning the bonded distance loss weight to enhance model performance.
Table 10: Comparison of ring distributions between molecules in the QM9 training set and those generated by our model.
3-Ring 4-Ring 5-Ring 6-Ring (7+)-Rings
QM9 train set 43.14% 39.20% 39.31% 12.72% 2.92%
Ours 41.69% 38.20% 36.37% 12.90% 3.13%
Ring Distribution. Following the evaluation protocol of [41], we computed the Percentage (%) of molecular modes in terms of ring distribution for our generated molecules and compared them against the QM9 training set. The results in Table 10 provide a quantitative assessment of how well our model preserves the structural diversity of the training dataset.
Figure 6: Additional t-SNE visualizations of UAE-3D’s latents under SE(3) augmentations. (a) 360-degree rotation along the z-axis. (b) same molecule under different SE(3) augmentations.
More Visualizations of Latent Space. In addition to Figure 4, we conducted more t-SNE visualizations of UAE-3D’s molecular latent representations to further examine its geometric sensitivity and consistency. In Figure 6(a), we apply 20-step rotations along the same axis (z-axis) to different molecules, completing a full 360-degree rotation. Remarkably, the latent trajectories of all molecule form loops and return to their original positions, reflecting the geometric sensitivity rotational consistency in the learned representations. In Figure 6(b), each color now corresponds to a different SE(3) augmentation applied to the same molecule, rather than to different molecules as in previous plots. We observe clear separation among representations generated by translations and rotations along different axes (x/y/z), indicating that UAE-3D effectively distinguishes between types of geometric transformations in its latent space. These results further validate the structured and disentangled nature of UAE-3D’s latent geometry.
B.2 More Visualizations of Generated Molecules
To further demonstrate the chemical validity and structural diversity of molecules generated by UDM-3D, we provide extended visual comparisons across the GEOM-Drugs and QM9 datasets in Figure 7 and Figure 8.
Figure 7: Visualization of random samples generated by UDM-3D on GEOM-Drugs.
Figure 8: More visualization of random samples generated by UDM-3D on QM9.
Appendix C More Related Works